OpenStudy (anonymous):

find limit then use l'hopital's rule to check it. lim of x approaches 2 for f(x)=[(x^2)-4]/[(x^2)+2x-8]

OpenStudy (anonymous):

answer is 2/3? says the book

OpenStudy (anonymous):

it's clear that when we subtitute dierctly by x=2 it will be undtermiend form , so using l'hopital rule ===> lim (2x/(2x-2) = 4/(4-2) =4/2 =2

OpenStudy (anonymous):

sorry it's lim (2x/(2x-2) = 4/(4+2) =4/6 =2/3

OpenStudy (anonymous):

I missed the positive sign before

OpenStudy (anonymous):

so just find the deravative of the numerator / derivative of denomerator , then subtitute for x=2 , you get the value

OpenStudy (anonymous):

hey thank you raheen, you've been such a great help. :)

OpenStudy (anonymous):

you are welcom, any time.